# Neural Network Diffusion

### v1

@@ 4 Is P-diff Only Memorizing?
(a) Similarity comparisons of original and p-diff models.
(b) Similarity comparisons of fine-tuned, noise-added, and p-diff models.
(c) t-SNE of the latent representations of original, p-diff, and adding noise.
Figure 4: The similarity represents the Intersection of Union (IoU) over wrong predictions between/among two models (a) shows the comparisons in four cases: similarity among original models and p-diff models, similarity between original and p-diff models, and the maximum similarity (nearest neighbor) between original and p-diff models. (b) displays the accuracy and max similarity of fine-tuned, noise-added, and p-diff models. All the maximum similarities are calculated with the original models. (c) presents the t-SNE (Van der Maaten et al., 2008) of latent representations of the original models, p-diff models, and adding noise operation.
@@
We conduct experiments on CIFAR-100 (Krizhevsky et al., 2009) with ResNet-18 (He et al., 2016) under the default setting, i.e. only generating the parameters of the last two batch normalization layers. We measure the similarity between the two models by calculating the Intersection over Union (IoU) on their wrong predictions. The IoU can be formulated as follows,
@@
We evaluate the similarity between the original and p-diff models. For each model, we obtain its similarity by averaging the IoUs with other models. We introduce four comparisons: 1) similarity among original models; 2) similarity among p-diff models; 3) similarity between original and p-diff models; and 4) max similarity (nearest neighbor) between original and p-diff models. We calculate the IoUs for all models in the above four comparisons and report their averaged values in Fig. 4(a).
@@
Based on the results in Fig. 4(b), we find that the performances of fine-tuned and noise-added models are hard to outperform the original models. Besides, the similarities between fine-tuned or noise-added and original models are very high, which indicates these two operations can not obtain novel but high-performing models. However, our generated models achieve diverse similarities and superior performances compared to the original models.
@@
In addition to predictions, we assess the distributions of latent representations for the original and generated models using t-SNE (Van der Maaten et al., 2008). To identify the differences between our approach and the operation of adding noise to the latent representations of original models, we also include the adding noise operation as a comparison in Fig. 4(c). The added noise is random Gaussian noise with an amplitude of 0.1. One can find that p-diff can generate novel latent representations while adding noise just makes interpolation around the latent representations of original models.
(a) Visualization of parameter trajectories of p-diff.
[…]
@@ 5 Related Work
HyperNet (Ha et al., 2017) dynamically generates the weights of a model with variable architecture. Smash (Brock et al., 2018b) introduces a flexible scheme based on memory read-writes that can define a diverse range of architectures.  (Peebles et al., 2023) collect 23 million checkpoints and train a conditional generator via a transformer-based diffusion model. MetaDiff (Zhang & Yu, 2023) introduces a diffusion-based meta-learning method for few-shot learning, where a layer is replaced by a diffusion U-Net (Ronneberger et al., 2015). HyperDiffusion (Erkoç et al., 2023) directly utilizes a diffusion model on MLPs to generate new neural implicit fields. Different from them, we analyze the intrinsic differences between images and parameters and design corresponding modules to learn the distributions of the high-performing parameters.
@@
Our approach could be viewed as learning a prior over network parameters, represented by the trained diffusion model. Learning parameter priors for neural networks has been studied in classical literature. Stochastic neural networks (SNNs) (Sompolinsky et al., 1988; Bottou et al., 1991; Wong, 1991; Schmidt et al., 1992; Murata et al., 1994) also learn such priors by introducing randomness to improve the robustness and generalization of neural networks. The Bayesian neural networks (Neal, 2012; Kingma & Welling, 2013; Rezende et al., 2014; Kingma et al., 2015; Gal & Ghahramani, 2016) aims to model a probability distribution over neural networks to mitigate overfitting, learn from small datasets, and asses the uncertainty of model predictions. (Graves, 2011) propose an easily implementable stochastic variational method as a practical approximation to Bayesian inference for neural networks. They introduce a heuristic pruner to reduce the number of network weights, resulting in improved generalization. (Welling & Teh, 2011) combine Langevin dynamics with SGD to incorporate a Gaussian prior into the gradient. This transforms SGD optimization into a sampling process. Bayes by Backprop (Blundell et al., 2015) learns a probability distribution prior over the weights of a neural network. These methods mostly operate in small-scale settings, while p-diff shows its effectiveness in real-world architectures.
@@ 6 Discussion and Conclusion
Neural networks have several popular learning paradigms, such as supervised learning (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014; He et al., 2016; Dosovitskiy et al., 2020), self-supervised learning (Devlin et al., 2018; Brown et al., 2020; He et al., 2020, 2022), and more. In this study, we observe that diffusion models can be employed to generate high-performing and novel neural network parameters, demonstrating their superiority. Using diffusion steps for neural network parameter updates shows a potentially novel paradigm in deep learning.
However, we acknowledge that images/videos and parameters are signals of different natures, and this distinction must be handled with care. Additionally, even though diffusion models have achieved considerable success in image/video generation, their application to parameters remains relatively underexplored. These pose a series of challenges for neural network diffusion. We propose an initial approach to address some of these challenges. Nevertheless, there are still unresolved challenges, including memory constraints for generating the entire parameters of large architectures, the efficiency of structure designs, and performance stability.
@@ Acknowledgments.
We thank Kaiming He, Dianbo Liu, Mingjia Shi, Zheng Zhu, Bo Zhao, Jiawei Liu, Yong Liu, Ziheng Qin, Zangwei Zheng, Yifan Zhang, Xiangyu Peng, Hongyan Chang, David Yin, Dave Zhenyu Chen, Ahmad Sajedi, and George Cazenavette for valuable discussions and feedbacks.
Appendix A Experimental Settings

### v2

@@ 4 Is P-diff Only Memorizing?
We conduct experiments on CIFAR-100 [36] with ResNet-18 [25] under the default setting, i.e. only generating the parameters of the last two batch normalization layers. We measure the similarity between the two models by calculating the Intersection over Union (IoU) on their wrong predictions. The IoU can be formulated as follows,
@@
We evaluate the similarity between the original and p-diff models. For each model, we obtain its similarity by averaging the IoUs with other models. We introduce four comparisons: 1) similarity among original models; 2) similarity among p-diff models; 3) similarity between original and p-diff models; and 4) max similarity (nearest neighbor) between original and p-diff models. We calculate the IoUs for all models in the above four comparisons and report their averaged values in Fig. 3(f).
@@
Based on the results in Fig. 3(g), we find that the performances of fine-tuned and noise-added models are hard to outperform the original models. Besides, the similarities between fine-tuned or noise-added and original models are very high, which indicates these two operations can not obtain novel but high-performing models. However, our generated models achieve diverse similarities and superior performances compared to the original models.
@@
In addition to predictions, we assess the distributions of latent representations for the original and generated models using t-SNE [71]. To identify the differences between our approach and the operation of adding noise to the latent representations of original models, we also include the adding noise operation as a comparison in Fig. 3(h). The added noise is random Gaussian noise with an amplitude of 0.1. One can find that p-diff can generate novel latent representations while adding noise just makes interpolation around the latent representations of original models.
(a) Parameter trajectories of p-diff.
(b) IoUs of generated models (Acc.$\geq$76.5%).
@@ 5 Related Work
HyperNet [22] dynamically generates the weights of a model with variable architecture. Smash [7] introduces a flexible scheme based on memory read-writes that can define a diverse range of architectures. G.pt [52] collects 23 million checkpoints (including accuracy, loss, and error) and trains a conditional generator using a transformer-based diffusion model. This model takes the current parameters as input and uses the target loss or error as a condition to generate new parameters. However, it is challenging to achieve results comparable to those of their training data (models). MetaDiff [76] introduces a diffusion-based meta-learning method for few-shot learning, where a layer is replaced by a diffusion U-Net [60]. It does not have the capability to generate high-performing neural network weights from random noise. Diffusion-SDF [9] proposes a diffusion model for shape completion, single-view reconstruction, and reconstruction of real-scanned point clouds. They still use the diffusion model for visual generation, not for parameter generation. HyperDiffusion [18] uses a diffusion model on MLPs to generate new neural implicit fields (a representation for 3D shape). DWSNets [47] introduces a symmetry-based approach for designing neural architectures that operate in deep weight spaces. NeRN [1] shows that neural representations can be used to directly represent the weights of a pre-trained convolutional neural network. Different from them, we analyze the intrinsic differences between images and parameters and design corresponding modules to learn the distributions of the high-performing parameters. More differences analysis are included in Appendix A.1 A.2 A.3.
@@
Our approach could be viewed as learning a prior over network parameters, represented by the trained diffusion model. Learning parameter priors for neural networks has been studied in classical literature. Stochastic neural networks (SNNs) [66, 5, 74, 62, 46] also learn such priors by introducing randomness to improve the robustness and generalization of neural networks. The Bayesian neural networks [48, 32, 58, 33, 20] aim to model a probability distribution over neural networks to mitigate overfitting, learn from small datasets, and estimate the uncertainty of model predictions. [21] propose an easily implementable stochastic variational method as a practical approximation to Bayesian inference for neural networks. They introduce a heuristic pruner to reduce the number of network weights, resulting in improved generalization. [72] combine Langevin dynamics with SGD to incorporate a Gaussian prior into the gradient. This transforms SGD optimization into a sampling process. Bayes by Backprop [3] learns a probability distribution prior over the weights of a neural network. These methods mostly operate in small-scale settings, while p-diff shows its effectiveness in real-world architectures.
@@ 6 Discussion and Conclusion
Neural networks have several popular learning paradigms, such as supervised learning [37, 64, 25, 17], self-supervised learning [14, 8, 24, 23], and more. In this study, we observe that diffusion models can be employed to generate high-performing and novel neural network parameters, demonstrating their superiority. Using diffusion steps for neural network parameter updates shows a potentially novel paradigm in deep learning.
However, we acknowledge that images/videos and parameters are signals of different natures, and this distinction must be handled with care. Additionally, even though diffusion models have achieved considerable success in image/video generation, their applications to parameters remain relatively underexplored. These pose a series of challenges for neural network parameter diffusion. We propose an initial approach to address some of these challenges. Nevertheless, there are still unresolved challenges, including memory constraints for generating the entire parameters of large architectures, the efficiency of structure designs, and performance stability.
@@ Acknowledgments.
We thank Kaiming He, Dianbo Liu, Mingjia Shi, Zheng Zhu, Bo Zhao, Jiawei Liu, Yong Liu, Ziheng Qin, Zangwei Zheng, Yifan Zhang, Xiangyu Peng, Hongyan Chang, David Yin, Dave Zhenyu Chen, Ahmad Sajedi, and George Cazenavette for valuable discussions and feedbacks. This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008).
Appendix A Frequent Asked Questions
We list the potential frequent asked questions and the point-to-point answers as follows:
A.1 Differences between G.pt and p-diff
G.pt [52] uses diffusion process as an optimizer (from starting parameter $\theta$ to future parameter $\theta^{*}$ defined by their paper), so they train 23 million checkpoints (2 layer MLP with 7960 parameters and CNNs with 5370 parameters) and save their loss and error.
In G.pt, the input consists of $\theta$, noised $\theta^{*}$, and loss and error of $\theta^{*}$ as conditions. However, we directly generate high-performing parameters from noise.
In Fig 4 of G.pt paper, it can not achieve similar results as Adam. For example, the error of G.pt is 50 while Adam is 40. P-diff can achieve similar or even better performances than models optimized by SGD. We show the detailed comparisons between G.pt and our approach in the following Tab. 5.
items G.pt p-diff
target optimizer, as an alternative for SGD and Adam generate high-performing parameters from noise
data preparation saving loss, error, and checkpoints during training saving 200 converged checkpoints
input starting checkpoints, noise, loss and error of starting and future checkpoints random noise
performance much worse than Adam, 10 percent drops can be found in Fig. 4 right one similar or even better than the best results of input models
implemented architectures MLP (2-layer, 10 hidden units); Conv + FC BN, Conv, FC, and entire parameters that include BN, Conv, and FC
generated parameter number less than 10, 000 2000 to 150, 000 (limited by the memory of hardware)
evaluating datasets MNIST, CIFAR-10 MNIST, CIFAR-10/100, STL-10, Flowers, Pets, F-101, ImageNet, COCO
evaluating tasks classification classification, segmentation, detection, visual generation
Table 5: Main differences between G.pt and p-diff.
Table 6: P-diff can precisely generate models with predefined performances.
accuracy range, i.e., condition p-diff (best)
97.34±$\pm 0.2$ 97.33
98.06±$\pm 0.3$ 98.01
98.97±$\pm 0.2$ 98.93
99.40±$\pm 0.1$ 99.67
Can we control the performance of the generated models?
G.pt [52] use the diffusion model as an optimizer. They train 23 million models and save the training loss and error as conditions. Therefore, in the inference phase, G.pt is expected to generate the model parameters with the input conditions. However, based on the results from Fig. 4 of their paper, G.pt can hardly achieve comparable results as the condition defined. There is a 10 percent drop of classification result in G.pt, indicating its poor ability to control the generated performance. To evaluate whether p-diff can accurately generate the models with defined conditions, we divide the SGD-trained models into several categories by the performances. As shown in Tab. 6, we set 4 accuracy ranges for the trained ConvNet-3 models. Then, we use one-hot vectors to encode these levels as conditions. We report the best generated results of different conditions in Tab. 6. Experimental results show that our method can precisely generate the models with different performances. This ability is crucial for real-world scenarios.
Conclusions.
  • G.pt is designed for an optimizer (23 million pretrained models), while p-diff aims to be a high-performing network parameter generator (200 pretrained models).
  • G.pt needs a starting parameter, error and loss of starting and target parameters, and noise. We only need noise.
  • G.pt can not perform as well as their training data, while p-diff performs similarly or even better than the training data.
  • G.pt only evaluates their effectiveness on classification, we show the effectiveness of classification, segmentation, detection, and visual generation.
A.2 MetaDiff vs p-diff
MetaDiff is not designed to generate high-performing parameters across an entire neural network but rather targets a specific part of it. According to their paper: "MetaDiff $\epsilon_{\theta}(\cdot)$ acts as a meta-optimizer, accepting the features and labels of all support samples $(u_{i},y_{i})\in S$ as inputs. It then learns target weights $w_{0}$ for the base learner $g_{w}(\cdot)$ from initial weights $w_{T}$ through a denoising process (refer to Fig.2(b) from their paper)." It replaces a neural network layer with diffusion steps. However, this methodology introduces several drawbacks:
1. It substantially increases the optimization cost during training.
2. There is a significant increase in the variability of testing outcomes, stemming from uncertainties around the parameters generated in the inner loop.
In conclusion, while MetaDiff presents an innovative approach, it does not have the capability to generate high-performing neural network weights from random noise.
A.3 Detailed discussions with more related works
We also discuss the following works in detail.
[A] [63]. Hyper-Representations as Generative Models: Sampling Unseen Neural Network Weights, 2022
[B] [79]. Hypertransformer: Model generation for supervised and semi-supervised few-shot learning, 2022
[C] [34]. Parameter prediction for unseen deep architectures, 2021
[D] [13]. Predicting parameters in deep learning, 2013
[E] [47]. Equivariant Architectures for Learning in Deep Weight Spaces, 2023
[F] [35]. Can We Scale Transformers to Predict Parameters of Diverse ImageNet Models? 2023
[A] vs p-diff.
[A] extends hyper-representations of model weights to sample new model weights from representation space. They build a model zoo that includes a lot of trained models. Then, an AE is trained to generate models. We use the diffusion model to learn the distribution of high-performing models and generate novel models from random noise. We also compare with [A] and find our p-diff performs better than  [63].
[B] vs p-diff.
[B] proposes HyperTransformer, which is used to generate models for supervised and semi-supervised tasks where labeled data are necessary. It relies on image features to generate small target CNN parameters, while p-diff generates neural network parameters from noise. The main comparison in HyperTransformer is other few-shot learning methods, e.g., MAML, and RFS. However, p-diff generates models mainly compare with SGD-trained models. HyperTransformer is designed for tasks where there are labels in the data and the neural network needs to be small CNN, while p-diff is designed for all neural networks regardless of data in tasks. Based on the above analysis, HyperTransformer is largely different from p-diff.
[C] vs p-diff.
[C] (i.e. GHN-2) improves Graph Hypernetworks (GNH) [77] by updating the parameters on various GHNs and introducing normalization on the predicted parameters. Meanwhile, they also introduce a dataset named DeepNets-1M to boost the training performance.
However, GHN-2 can hardly achieve comparable results as the SGD (or other optimizers)-trained models, so it is usually regarded as an initialization method. Besides, GHN-2 is designed for NAS (neural architecture search) training. Therefore, GHN-2 predicts parameters through the given architectures and the authors mainly focus on improving the training efficiency on unseen architectures. They want to reduce the training cost of unseen architectures in NAS tasks. They do not consider too much of the diversity of generated models.
[D] vs p-diff.
[D] finds that the weights of a trained network exist 95% redundancy. They only use 5% weights to predict the remaining weights. The paper also proposes that these remaining weights may not even be learned. In other words, by training only a small part (5%) of the original parameters, it is possible to achieve performance similar to or even exceeding the original network (can be regarded as a kind of regularization). 1. P-diff does not need the trained weights as inputs for predicting the other weights. 2. P-diff mainly aims at generating high-performing but novel models, not consider too much about the redundancy of parameters.
[E] vs p-diff.
[E] introduces a symmetry-based approach for designing neural architectures that operate in deep weight spaces. The authors focus on what neural architectures can effectively learn and process neural models that are represented as sequences of weights and biases. However, p-diff does not focus on designing neural architectures but on generating neural network parameters from random noise.
[F] vs p-diff.
[F] (i.e. GHN-3) is also an initialization method and can not achieve similar results as SGD (or other optimizers)-trained models. Its motivation is to reduce the cost of pretraining the ImageNet. The authors introduce transformer to improve the efficiency and scalability of GHNs. GHN-3 + finetuning (less epochs) can achieve comparable results as SGD-trained models. P-diff can directly achieve similar or even better results than SGD-trained models without additional finetuning. Both p-diff and GHN series (GHN, GHN-2, GHN-3) can provide initialization for network training.
A.4 Bottleneck of scaling up
We estimate the memory cost of applying AE and LDM in different numbers of generated parameter.
Setting and details:
1. We record the GPU’s memory cost in models with different parameter numbers.
2. There are two stages in training p-diff, AE, and diffusion.
3. We set the batch size as 200 (the same as our paper).
4. We report the memory cost using MB in two stages.
Findings and conclusions:
1. As shown in Tab. 7, the cost of memory increases very fast with the number of parameters.
2. We empirically find that the memory cost of generating 100K parameters raises to the upper bound of A100 40G.
3. In our entire parameter generation (Tab. 3(b)), the maximum number of generated parameters is 150K (MLP on CIFAR-100). We reduce the default batch size from 200 to 50. That increases the training time a lot, so hardware is really a bottleneck in scaling up our method.
Table 7: Bottleneck of scaling up.
parameter number (K) AE Diffusion
2 1475 1501
10 4473 4514
70 23377 23491
100 40380 40581
A.5 Minor over baseline
Our model is trained from SGD pretrained models, so we target to achieve similar performance as SGD. More importantly, we focus on exploring whether the generated models only memorize the training data. Like visual generation, several works [78, 70] utilize diffusion as a data augmentation to boost the original results (as diffusion can generate something new). Here, our experiments demonstrate that p-diff can also generate some novel but high-performing models.
A.6 Main change seems to be having latent diffusion?
We show the differences between G.pt and p-diff in Tab. 5. In technical, our contributions can be summarized as follows:
1. Considering the difference between visual signals and neural network weights, we design 1D CNNs in the autoencoder and LDM.
2. To improve the performance stability of generated models, we introduce noise augmentation on both input and latent representations. It is worth noting that this operation is different from VAE.
3. We propose a metric for evaluating the difference between the two models.
4. We comprehensively analyze the characteristics of the generated models, including showing the differences between the original and generated models.
In the fresh views aspect:
1. P-diff can provide good initialization to speed up the training. (results can be found in Tab. 9)
2. P-diff can be a customized parameter generator. We evaluate it possibility in Tab. 6
3. Several long-way-to-go targets: can we use AI to generate AI? Can we use this technology to achieve a ModelGPT? People just need to tell the computer one sentence, ModelGPT provides the model. We believe our research can inspire more and more people to explore this area.
A.7 Other baselines
We also compare our approach with hyper-representations method [63]. We report the comparison results on MNIST, SVHN, CIFAR-10, and STL-10 datasets. To make a fair comparison, we also implement our approach using the three-layer convolutional neural network defined in hyper-representations papers [63]. As illustrated in Tab. 8, we can achieve lossless or even improved results than the original models, but the hyper-representations method can not obtain lossless results in most cases.
Table 8: Comparisons between p-diff and hyper-representations.
dataset original hyper-representations p-diff improvement
MNIST 93.7 93.8 93.9 $\uparrow$ 0.1
SVHN 78.3 78.4 78.6 $\uparrow$ 0.2
CIFAR-10 49.4 49.1 49.5 $\uparrow$ 0.4
STL-10 39.4 39.3 39.6 $\uparrow$ 0.3
A.8 Can p-diff be applied for parameter initialization?
To answer this question, we design experiments as follows. We consider two cases, the first one is partial parameter generation, and the second one is entire parameter generation. We implement our approach with ResNet-18 and ConvNet-3 on CIFAR-10 and CIFAR-100, respectively. We report the epoch number at which the model converges for evaluation. As shown in Tab. 9, our method speeds up the original by at least 5 times, which indicates the benefit of our method.
Table 9: P-diff performs good for parameter initialization.
dataset architecture generation mode original p-diff speed up ratio
CIFAR-10 ResNet-18 partial 11 epochs 2 epochs $\uparrow$ 5.50 $\times$
CIFAR-100 ResNet-18 partial 21 epochs 4 epochs $\uparrow$ 5.25 $\times$
CIFAR-10 ConvNet-3 entire 100 epochs 11 epochs $\uparrow$ 9.09 $\times$
CIFAR-100 ConvNet-3 entire 100 epochs 17 epochs $\uparrow$ 7.69 $\times$
A.9 Details of the parameter size in this work
The parameter number in p-diff experiments is from 2000 to 150, 000 (limited by the memory of hardware). We list the parameter number of our approach in the following table.
Table 10: Parameter numbers in different architectures in our experiments.
dataset architecture generation mode parameter number
CIFAR-100 ResNet-18 BN partial 2048
CIFAR-100 ResNet-18 shortcut partial 8192
CIFAR-100 ResNet-18 linear partial 51200
CIFAR-10 ConvNet-3 entire 24714
CIFAR-10 MLP entire 39760
CIFAR-100 ConvNet-3 entire 70884
CIFAR-100 MLP entire 155135
A.10 Why access to training set when selecting generated model?
As the normal SGD training process can only access the training set, we thus select the models based on their results on the training dataset for a fair comparison. Meanwhile, the experimental results in Fig. 6(a) indicate that p-diff exhibits a high level of consistency between the training and validation sets. We use item avg. in Tab. 3(e) 5(c) 5(e) 5(h) 5(k) to report the average performance of the 100 novel parameters generated when evaluated directly on the test set.
A.11 In-dataset fine-tuning comparison with hyper-representations [63].
We finetune ResNet-18 normalization layers with CIFAR-10 and save finetuned weights at different epochs. We train it for 10 epochs. Then, we use weights at different epochs to train p-diff and autoencoder with 1D CNNs. Finally, finetuning the weights to the 10 epoch (same as training). We report finetune performance between original, AE + Kernel Density Estimation (KDE) , and p-diff. As shown in Tab. 11, the models initialized with generated weights learn faster and achieve better results.
Table 11: Results of in-dataset fine-tuning. ‘-’ denotes empty (10 + 0 epoch fine-tuning).
Method/Epoch 0 1 5 10
original 13.7 40.7 90.4 92.2
AE+KDE [63] 91.2 91.9 92.3 -
p-diff 92.2 92.2 92.5 -
A.12 How about using Kernel Density Estimation(KDE) on top of the autoencoder?
We implement KDE on top of the autoencoder with our default setting that CIFAR-100 dataset and ResNet-18 backbone. We report the best, average, and median accuracy of generated models by p-diff, VAE and autoencoder with KDE sampling. Besides, we investigate the diversity of models generated by different methods. We calculate the similarity of 100 models generated by each method that performs similarly to the original models. The minimal similarity is shown in Tab. 12. AE+KDE performs better than VAE. P-diff achieves the best results on best, average, and median accuracy. Based on the similarity, p-diff generated models are more diverse than VAE and AE+KDE generated.
Table 12: Performance and similarity comparisons among original, VAE, AE+KDE, and p-diff.
method best avg. med. minimal similarity
original 76.6 - - 1.00
VAE 76.7 62.7 70.8 0.92
AE+KDE 76.7 70.3 73.4 0.93
p-diff 76.9 72.4 75.6 0.85
Appendix B Experimental Settings